# Post-Incorporating Code Structural Knowledge into Pretrained Models via ICL for Code Translation

### v1

@@ Abstract
Code translation migrates codebases across programming languages. Recently, large language models (LLMs) have achieved significant advancements in software mining. However, handling the syntactic structure of source code remains a challenge. Classic syntax-aware methods depend on intricate model architectures and loss functions, rendering their integration into LLM training resource-intensive. This paper employs in-context learning (ICL), which directly integrates task exemplars into the input context, to post-incorporate code structural knowledge into pre-trained LLMs. We revisit exemplar selection in ICL from an information-theoretic perspective, proposing that list-wise selection based on information coverage is more precise and general objective than traditional methods based on combine similarity and diversity. To address the challenges of quantifying information coverage, we introduce a surrogate measure, Coverage of Abstract Syntax Tree (CAST), measuring maximum subtree coverage between ASTs of test source code and exemplars. Furthermore, we formulate the NP-hard CAST maximization for exemplar selection and prove that it is a standard submodular maximization problem. Therefore, we propose a greedy algorithm for CAST submodular maximization, which theoretically guarantees a $(1-1/e)$-approximate solution in polynomial time complexity. Our method is the first training-free and model-agnostic approach to post-incorporate code structural knowledge into existing LLMs at test time. Experimental results show that our method significantly improves LLMs performance in code translation and reveals two meaningful insights: 1) Code structural knowledge can be effectively post-incorporated into pre-trained LLMs during inference, despite being overlooked during training; 2) Scaling up model size or training data does not lead to the emergence of code structural knowledge, underscoring the necessity of explicitly considering code syntactic structure.
Index Terms:
Code Translation, Syntactic Validity, Large Language Model, In-context Learning.
publicationid: pubid: $0$–$00$ 20xx IEEE
@@ I Introduction
Recently, large language models (LLMs) have demonstrated exceptional performance across various software mining tasks [1, 2, 3, 4]. However, effectively handling the syntactic structure of source code remains a significant challenge. Unlike the grammatical structure of natural language, the syntactic structure of source code is stricter and more complex, as it directly reflects program behavior. Previous studies have shown that explicitly incorporating code syntactic knowledge can significantly enhance performance in software mining tasks [5, 6, 7, 8, 9, 10, 11, 12]. However, traditional methods often depend on complex model architectures and customized loss functions, making their integration with LLMs both time-consuming and resource-intensive, even infeasible for proprietary models.
In-context learning (ICL) provides a effective and efficient solution of integrating new domain-specific knowledge into off-the-shelf LLMs without additional training by directly incorporating task exemplars into the input context [13]. For instance, in Python-to-C++ code translation, ICL enhances LLMs performance by incorporating a few paired exemplars (e.g., Python-to-C++ translation pairs) into the input prompt. This training-free approach offers flexibility across diverse tasks, avoids catastrophic forgetting, and ensures efficient resource utilization. Selecting appropriate contextual exemplars is critical for ICL. However, research on exemplar selection strategies for ICL in code translation is still limited.
@@
To address these challenges, we register a unique fingerprint to each AST subtree. Fingerprints for all subtrees in an AST are efficiently extracted via a single post-order traversal, and the exemplar database can be pre-processed efficiently. Next, we construct a co-occurrence matrix $M$ between the test source code and candidate exemplars, where $M_{ij}$ denotes whether the $j$-th AST subtree of the source code exists in the $i$-th candidate exemplar. Consequently, exemplar selection can be formulated as selecting rows from the co-occurrence matrix to maximize the $\ell_{1}$-norm of the OR vector across selected rows. We prove this is a standard submodular maximization problem, which has been widely studied in subset selection area. Although this is a general NP-hard problem, we can employ a greedy algorithm that starts with an empty selection set and iteratively adds elements to maximize the marginal gain of CAST at each step until the size limit $k$ is reached. This algorithm theoretically guarantees a $(1-1/e)$-approximate solution with polynomial time complexity, i.e., its performance is at least $(1-1/e)$ of the optimal solution.
@@
    Practical solution for CAST maximization: Although CAST maximization in exemplar selection is NP-hard, we prove it is a submodular maximization problem and propose a greedy algorithm that theoretically guarantees a $(1-1/e)$-approximate solution with practical polynomial time complexity. Experimental results show that CAST significantly improves the code translation performance of LLMs across various scales.
II Revisiting Exemplar Selection via
Information Theory for ICL in Code Translation
@@ III Proposed Method
To enhance code translation in LLMs via ICL, we propose CAST, a surrogate measure for information coverage, and present a practical solution for NP-hard CAST maximization with exemplar set selection. In Sec. III-A, we define CAST, a list-wise $\mathrm{Score}$ function based on the code’s AST. CAST accounts for both the relevance of exemplars to the source code and the compositional relationships among multiple selected exemplars. In Sec. III-B, we present a greedy algorithm for CAST maximization, proving that it guarantees a $(1-1/e)$-approximation for the NP-hard subset selection problem with polynomial time complexity. Finally, Sec. III-C provides an overview of the CAST framework.
@@ III-A Surrogate Scoring Function: CAST
[…]
@@ III-B Practical Solution: Greedy Submodular Maximization
Maximizing a non-negative, monotone, and submodular function in subset selection referred to submodular maximization. Although generally NP-hard, a greedy algorithm that iteratively maximizes the marginal gain offers a theoretical $(1-1/e)$-approximate solution in polynomial time. The surrogate CAST (Eq. 10) corresponds to the value function $f(S)=\left|\bigvee_{s\in S}M[s,:]\right|_{1}$ in the optimization problem (Eq. 12). This value function is non-negative, monotone, and submodular. We present a lemma to prove these properties:
@@
The $\ell_{1}$-norm of a binary vector $A\in\{0,1\}^{m}$ counts the number of elements equal to “$1$”:
@@
Theorem 4 ( $(1-1/e)$-approximation of greedy submodular maximization [26]).
@@ 1 Initialize the selected set: $S\leftarrow\emptyset$;
[…]
@@ 16       Update selected set: $S\leftarrow S\cup\{s\}$;
As shown in Theorem 4, our CAST maximization method in Eq. 24 and Algorithm 2 provides a $(1-1/e)$-approximation, ensuring sufficient effectiveness. For instance, if $k=3$ exemplars are selected, the approximation ratio reaches at least $1-(1-1/3)^{3}\approx 70.37\%$.
@@
In summary, constructing the co-occurrence matrix for CAST takes $\mathcal{O}(n\cdot|\Phi_{\text{test}}|)$, while the greedy CAST algorithm in Algorithm 2 has a polynomial time complexity of $\mathcal{O}(nk\cdot|\Phi_{\text{test}}|)$. Using distributed and parallel computation, this can be further reduced to $\mathcal{O}(k\cdot|\Phi_{\text{test}}|)$. Thus, our approach ensures both theoretical effectiveness with a near-optimal $(1-1/e)$-approximation and practical efficiency.
@@ IV-A Dataset and Database
We use the dataset from [28], containing 948 parallel functions in Python, Java, and C++, widely used in code translation [28, 29, 30]. Of these, 568 include unit tests for at least one programming language: 464 for Python, 482 for Java, and 467 for C++. Following the previous work [31], our experiments focus on this subset. ICL requires a separate database. For each pre-trained LLM, we use it to translate the entire test dataset and verify the predictions using test input cases. Samples that pass all test input cases are added to the database for this LLM. Models are evaluated on the remaining data, and final metrics are computed using the entire dataset, treating the database samples as exactly correct.
@@ IV-B Baselines
TABLE I: Results of pre-trained models, including LLMs and code translation models, with fixed exemplar strategy.
@@ IV-B1 Baseline Code Translation Models
  • GPT-4 [38]: Advanced generative AI models developed by OpenAI. Although not explicitly trained for code generation, they demonstrate notable performance in this domain. Their effectiveness in handling code generation tasks is primarily attributed to their massive scale.
Moreover, three state-of-the-art learning-based code translation models are served as baselines.
@@ IV-B2 Baseline ICL Exemplar Selection Strategies
We evaluate several baseline ICL exemplar selection strategies to compare with our proposed CAST variants. These strategies encompass simple heuristic methods, similarity-based approaches, and diversity-oriented techniques, as detailed below:
  • Fixed Exemplars (Fixed): A static set of exemplars is pre-selected and consistently applied to all test samples.
  • Random Sampling (Random): Exemplars are randomly selected from the database for each test sample, serving as a baseline for comparison without any guided selection.
  • Similarity with Levenshtein Distance (LD): Exemplars are selected via lexical similarity to the test source code, measured by the Levenshtein distance.
  • Similarity on CodeBERT Embedding (CodeBERT): Exemplars are retrieved using cosine similarity scores between CodeBERT [39] embeddings of the test source code and candidate samples.
  • Diversity with Clustering (Diversity): Candidate samples are grouped into $k$ clusters by $k$-means based on the CodeBERT [39] embeddings, and the most similar sample from each cluster is selected to ensure diversity.
@@ IV-D Empirical Results and Analysis
[…]
@@
question-circle RQ1: How do SOTA LLMs of various scales perform in the code translation task?
@@
In this research question, we evaluate SOTA LLMs across various scales for code translation. As shown in Tab. I, the results reveal several meaningful observations: 1) The LLMs demonstrate impressive performance in the code translation task, highlighting their potential for practical application. Specifically, the proprietary LLMs, such as GPT-40 and GPT-4o-mini, achieve an average CA of 66.37% and 69.72%, respectively, indicating strong potential for LLM-based code translation. 2) A high EM score indicates that the translated code is identical to the ground truth, often used as a metric to assess the extent of training data leakage. Compared to learning-based methods, which ensure no data leakage during training, the EM scores of LLMs are comparable, indicating minimal data leakage. This shows that LLMs can effectively understand the task through prompts and produce semantically correct translated code. 3) Most LLMs outperform learning-based models. Within the same model series, larger models, such as the Gemma2 series, perform better than smaller ones, indicating that scaling up parameters enhances the code translation capabilities of LLMs.
[…]
@@
[…]
question-circle RQ2: How does CAST perform with SOTA LLMs compared to different ICL exemplar selection strategies?
Referring to the results in Tab. I, we choose three open-source LLMs as primary baselines: Qwen2-7B, the SOTA LLM under 10B parameters; Gemma2-27B, the SOTA LLM under 40B parameters; and Qwen2.5-72B, the SOTA LLM exceeding 70B parameters. CAST-F is used to select exemplars for ICL, and its performance is compared against other exemplar selection strategies. Results in Tab. II demonstrate the effectiveness of CAST-F across LLMs with various parameter scales in the 5-shot code translation task. There are several key observations: 1) CAST-F consistently outperforms all other ICL exemplar selection strategies across LLMs and translation tasks. Specifically, it surpasses the second-best strategy, Diversity, by 1.6% for Qwen2-7B, 1.9% for Gemma2-27B, and 1.14% for Qwen2.5-72B. This shows that information coverage is a more precise and general measure for ICL, and CAST-F effectively integrates code structural knowledge into LLMs to enhance code translation. 2) For Gemma2-27B, exemplar selection using LD, CodeBERT, and Diversity performs worse than random sampling. This highlights the importance of effective exemplar selection strategies in ICL. CAST-F, the only list-wise method that considers both information coverage and code structural knowledge, consistently achieves superior results compared to other strategies.
[…]
TABLE III: Performance of Qwen2-7B on code translation with various ICL exemplar selection strategies and numbers of exemplars.
@@
[…]
question-circle RQ3: How does ICL perform using different exemplar selection strategies as the number of exemplars increases?
Results in Tab. III and Fig. 3 show the performance of Qwen2-7B in code translation using various ICL exemplar selection strategies, with the number of exemplars increasing from $1$-shot to $20$-shots. Several key observations arise from this analysis: 1) CAST-F consistently achieves the highest average performance across all strategies, with an accuracy of 65.11% at $20$-shots, surpassing the second-best strategy, Diversity, by 2.02%. 2) Increasing the number of exemplars generally improves performance across most strategies. CAST-F demonstrates superior scalability, particularly at higher shot counts, underscoring its robustness in improving code translation. 3) Compared to baselines, CAST-F consistently delivers incremental gains by leveraging information coverage and code structural knowledge. 4) The Diversity strategy ranks second, performing well by considering both point-wise similarity and diversity, but it lacks the ability to account for code structural knowledge or perform selection in a compositional manner. 5) CAST-F achieves the highest accuracy across challenging language pairs, such as Python to Java (58.10%) and Java to Python (69.37%), demonstrating its effectiveness in translating between syntactically diverse languages. 6) A law of diminishing marginal utility is evident in Tab. III and Fig. 3 as the number of exemplars increases. While performance improves significantly with the first few exemplars, additional gains become minimal beyond $5$-shots due to redundancy or information saturation. However, Qwen2-7B can remit this issue effectively by considering information coverage, as a more precise and general measure.
[…]
@@
question-circle RQ4: How does CAST-A perform when dynamically determining the selection size?
[…]
question-circle RQ5: How does the CAST convergence of CAST-F compare with that of LD?
@@
[…]
TABLE IV: Results of the gain of CAST-F among different sizes of LLMs, consisting of 2B, 9B, and 27B of Gemma2.
@@
[…]
question-circle RQ6: Does scaling up the model size automatically lead to sufficient code structural knowledge?
As shown in Tab. II, although translation performance improves with larger models, the improvement from explicitly incorporating the code structural knowledge by CAST-F is much more substantial. Additionally, for a fairer comparison, we evaluate CAST-F with $5$-shots on Gemma2-(2B/9B/27B), all of which are from the same family. The results, shown in Tab. IV, indicate that the improvement from CAST-F is not significantly related to model scale. This confirms that the additional code structural knowledge from CAST-F is independent of model scale. Therefore, simply scaling up the model or training data does not result in the sufficient emergence of code structural knowledge, highlighting the necessity of explicitly incorporating code syntax.
[…]
@@
[…]
question-circle RQ7: What is the hyperparameter sensitivity?

### v2

@@ Abstract
Code translation migrates codebases across programming languages. Recently, large language models (LLMs) have achieved significant advancements in software mining. However, handling the syntactic structure of source code remains a challenge. Classic syntax-aware methods depend on intricate model architectures and loss functions, rendering their integration into LLM training resource-intensive. This paper employs in-context learning (ICL), which directly integrates task exemplars into the input context, to post-incorporate code structural knowledge into pre-trained LLMs. We revisit exemplar selection in ICL from an information-theoretic perspective, proposing that list-wise selection based on information coverage is more precise and general objective than traditional methods based on combine similarity and diversity. To address the challenges of quantifying information coverage, we introduce a surrogate measure, Coverage of Abstract Syntax Tree (CAST), measuring maximum subtree coverage between ASTs of test source code and exemplars. Furthermore, we formulate the NP-hard CAST maximization for exemplar selection and prove that it is a standard submodular maximization problem. Therefore, we propose a greedy algorithm for CAST submodular maximization, which theoretically guarantees a (11/e)(1-1/e)-approximate solution in polynomial time complexity. Our method is the first training-free and model-agnostic approach to post-incorporate code structural knowledge into existing LLMs at test time. Experimental results show that our method significantly improves LLMs performance in code translation and reveals two meaningful insights: 1) Code structural knowledge can be effectively post-incorporated into pre-trained LLMs during inference, despite being overlooked during training; 2) Scaling up model size or training data does not lead to the emergence of code structural knowledge, underscoring the necessity of explicitly considering code syntactic structure.
@@ I Introduction
Recently, large language models (LLMs) have demonstrated exceptional performance across various software mining tasks [1, 2, 3, 4, 5, 6, 7]. However, effectively handling the syntactic structure of source code remains a significant challenge. Unlike the grammatical structure of natural language, the syntactic structure of source code is stricter and more complex, as it directly reflects program behavior. Previous studies have shown that explicitly incorporating code syntactic knowledge can significantly enhance performance in software mining tasks [8, 9, 10, 11, 12, 13, 14, 15]. However, traditional methods often depend on complex model architectures and customized loss functions, making their integration with LLMs both time-consuming and resource-intensive, even infeasible for proprietary models.
In-context learning (ICL) provides an effective and efficient solution of integrating new domain-specific knowledge into off-the-shelf LLMs without additional training by directly incorporating task exemplars into the input context [16]. For instance, in Python-to-C++ code translation, ICL enhances LLMs performance by incorporating a few paired exemplars (e.g., Python-to-C++ translation pairs) into the input prompt. This training-free approach offers flexibility across diverse tasks, avoids catastrophic forgetting, and ensures efficient resource utilization. Selecting appropriate contextual exemplars is critical for ICL. However, research on exemplar selection strategies for ICL in code translation is still limited.
@@
To address these challenges, we register a unique fingerprint to each AST subtree. Fingerprints for all subtrees in an AST are efficiently extracted via a single post-order traversal, and the exemplar database can be pre-processed efficiently. Next, we construct a co-occurrence matrix $M$ between the test source code and candidate exemplars, where $M_{ij}$ denotes whether the $j$-th AST subtree of the source code exists in the $i$-th candidate exemplar. Consequently, exemplar selection can be formulated as selecting rows from the co-occurrence matrix to maximize the $\ell_{1}$-norm of the OR vector across selected rows. We prove this is a standard submodular maximization problem, which has been widely studied in subset selection area. Although this is a general NP-hard problem, we can employ a greedy algorithm that starts with an empty selection set and iteratively adds elements to maximize the marginal gain of CAST at each step until the size limit $k$ is reached. This algorithm theoretically guarantees a (11/e)(1-1/e)-approximate solution with polynomial time complexity, i.e., its performance is at least (11/e)(1-1/e) of the optimal solution.
@@
    Practical solution for CAST maximization: Although CAST maximization in exemplar selection is NP-hard, we prove it is a submodular maximization problem and propose a greedy algorithm that theoretically guarantees a (11/e)(1-1/e)-approximate solution with practical polynomial time complexity. Experimental results show that CAST significantly improves the code translation performance of LLMs across various scales.
II Revisiting Exemplar Selection via Information Theory for ICL in Code Translation
@@ III Proposed Method
To enhance code translation in LLMs via ICL, we propose CAST, a surrogate measure for information coverage, and present a practical solution for NP-hard CAST maximization with exemplar set selection. In Sec. III-A, we define CAST, a list-wise $\mathrm{Score}$ function based on the code’s AST. CAST accounts for both the relevance of exemplars to the source code and the compositional relationships among multiple selected exemplars. In Sec. III-B, we present a greedy algorithm for CAST maximization, proving that it guarantees a (11/e)(1-1/e)-approximation for the NP-hard subset selection problem with polynomial time complexity. Finally, Sec. III-C provides an overview of the CAST framework.
@@ III-A Surrogate Scoring Function: CAST
$\text{node}_{1}$ $=\{\text{``type'': ``variable name'', ``value'': ``count''}\}\,$, (9)
$\text{node}_{2}$ $=\{\text{``type'': ``function name'', ``value'': ``sort'`}\}\,.$
@@ III-B Practical Solution: Greedy Submodular Maximization
Maximizing a non-negative, monotone, and submodular function in subset selection referred to submodular maximization. Although generally NP-hard, a greedy algorithm that iteratively maximizes the marginal gain offers a theoretical (11/e)(1-1/e)-approximate solution in polynomial time. The surrogate CAST (Eq. 10) corresponds to the value function $f(S)=\left|\bigvee_{s\in S}M[s,:]\right|_{1}$ in the optimization problem (Eq. 12). This value function is non-negative, monotone, and submodular. We present a lemma to prove these properties:
@@
The $\ell_{1}$-norm of a binary vector $A\in\{0,1\}^{m}$ counts the number of elements equal to “11”:
@@
Theorem 4 ( (11/e)(1-1/e)-approximation of greedy submodular maximization [29]).
@@ 4 $s\leftarrow-1$; $\text{max\_gain}\leftarrow-1$;
7       if $g>\text{max\_gain}$ then
8          $\text{max\_gain}\leftarrow g$;
@@ 16   Update selected set: $S\leftarrow S\cup\{s\}$;
As shown in Theorem 4, our CAST maximization method in Eq. 24 and Algorithm 2 provides a (11/e)(1-1/e)-approximation, ensuring sufficient effectiveness. For instance, if $k=3$ exemplars are selected, the approximation ratio reaches at least $1-(1-1/3)^{3}\approx 70.37\%$.
@@
In summary, constructing the co-occurrence matrix for CAST takes $\mathcal{O}(n\cdot|\Phi_{\text{test}}|)$, while the greedy CAST algorithm in Algorithm 2 has a polynomial time complexity of $\mathcal{O}(nk\cdot|\Phi_{\text{test}}|)$. Using distributed and parallel computation, this can be further reduced to $\mathcal{O}(k\cdot|\Phi_{\text{test}}|)$. Thus, our approach ensures both theoretical effectiveness with a near-optimal (11/e)(1-1/e)-approximation and practical efficiency.
@@ IV-A Dataset and Database
We use the dataset from [31], containing 948 parallel functions in Python, Java, and C++, widely used in code translation [31, 32, 33]. Of these, 568 include unit tests for at least one programming language: 464 for Python, 482 for Java, and 467 for C++. Following the previous work [34], our experiments focus on this subset. To further validate our approach, we evaluate various methods on AVATAR [35], a more realistic benchmark constructed from competitive programming sites, online platforms, and open-source repositories. We also extend our evaluation beyond code translation to code summarization using benchmarks from CodeXGLUE [36]. ICL requires a separate database. For each pre-trained LLM, we use it to translate the entire test dataset and verify the predictions using test input cases. Samples that pass all test input cases are added to the database for this LLM. Models are evaluated on the remaining data, and final metrics are computed using the entire dataset, treating the database samples as exactly correct. The replication package is publicly available at https://doi.org/10.5281/zenodo.16920815.
@@ IV-B Baselines
TABLE I: Results of pre-trained models, including LLMs and code translation models, with one-shot fixed exemplar strategy.
@@ IV-B1 Baseline Code Translation Models
  • GPT-4o [43]: Advanced generative AI models developed by OpenAI. Although not explicitly trained for code generation, they demonstrate notable performance in this domain. Their effectiveness in handling code generation tasks is primarily attributed to their massive scale.
Moreover, three state-of-the-art learning-based code translation models serve as baselines.
@@ IV-B2 Baseline ICL Exemplar Selection Strategies
We evaluate several baseline ICL exemplar selection strategies to compare with our proposed CAST variants.
These strategies encompass simple heuristic methods, similarity-based approaches (e.g., LD, CodeBERT, and BM25), and diversity-oriented techniques (e.g., Diversity), which are essentially different implicit surrogates of Sec. II-B. Considering only similarity is equivalent to setting $\lambda=0$ in Sec. II-B. Each exemplar selection method can be defined as a specific instantiation of the general formulation Eq. 3.
Fixed Exemplars (Fixed): A static exemplar set is shared across all test samples:
$S^{*}=S_{\text{fixed}},\quad\forall X_{\text{test}}\,.$ (25)
Random Sampling (Random): $k$ exemplars are uniformly sampled from the candidate pool, independent of $X_{\text{test}}$:
$S^{*}\sim\text{Uniform}(\mathbb{N}_{n}),\quad|S^{*}|=k\,.$ (26)
Retrieval by Levenshtein Distance (LD): The score is defined as the negative sum of Levenshtein distances between each exemplar and the test sample:
$S^{*}=\arg\max_{S\subseteq\mathbb{N}_{n},\,|S|=k}-\sum_{i\in S}\text{LD}(X_{i},X_{\text{test}})\,.$ (27)
Similarly, the method Retrieval by AST Edit Distance (AST-ED) is defined based on the Tree Edit Distance [44] between each exemplar and the test sample, where only the type information of AST nodes is retained for distance calculation to improve generalization, aligning with the setting of CAST.
Retrieval by CodeBERT Similarity (CodeBERT): Each sample is encoded using a CodeBERT embedding function $f_{\text{CB}}(\cdot)$, and cosine similarity is used:
$S^{*}=\arg\max_{S\subseteq\mathbb{N}_{n},\,|S|=k}\sum_{i\in S}\cos\left(f_{\text{CB}}(X_{i}),f_{\text{CB}}(X_{\text{test}})\right)\,.$ (28)
Retrieval by BM25 (BM25): The score is the sum of BM25 relevance scores between each candidate and the test input:
$S^{*}=\arg\max_{S\subseteq\mathbb{N}_{n},\,|S|=k}\sum_{i\in S}\text{BM25}(X_{i},X_{\text{test}})\,.$ (29)
Diversity with Clustering (Diversity): The candidate pool is clustered into $k$ groups using CodeBERT embeddings, and the most similar sample in each cluster is selected:
$\mathbb{N}_{n}\xrightarrow[]{\text{k-means}(f_{\text{CB}}(X_{i}))}\{C_{1},\dots,C_{k}\}\,$, (30)
where $C_{j}$ represents the index set of the $j$-th cluster.
$S^{*}_{[i]}=\arg\max_{j\in C_{i}}\cos\left(f_{\text{CB}}(X_{j}),f_{\text{CB}}(X_{\text{test}})\right)\,$, (31)
for $i\in\{1,\dots,k\}$, and $S^{*}=S^{*}_{[1:k]}$.
@@ IV-D Empirical Results and Analysis
GPT-4o Random 72.01% 69.19% 73.42% 70.95% 69.01% 71.48% 71.01%
LD 72.36% 70.25% 73.59% 71.48% 69.54% 71.48% 71.45%
AST-ED 72.71% 70.42% 73.59% 71.30% 69.54% 71.48% 71.51%
BM25 72.36% 70.60% 73.06% 71.83% 69.19% 71.30% 71.39%
CodeBERT 73.06% 69.89% 73.94% 71.48% 69.37% 71.48% 71.54%
Diversity 72.36% 69.54% 73.59% 71.13% 69.37% 71.65% 71.27%
CAST-F 73.24% 70.77% 74.12% 72.01% 70.07% 71.65% 71.98%
TABLE III: Results of various ICL strategies for 5-shot translation on AVATAR [35] of Qwen2-7B.
Strategy Java to Python Python to Java
CA CodeBLEU BLEU WNG Syntax. Dataflow. CA CodeBLEU BLEU WNG Syntax. Dataflow.
Random 39.36% 20.06 1.56 2.29 36.39 35.40 46.40% 20.88 2.01 3.10 43.88 29.56
LD 40.56% 19.73 1.56 2.23 34.77 33.40 47.59% 21.30 1.90 2.90 42.48 29.96
BM25 46.59% 19.85 1.61 2.26 36.03 35.50 44.00% 21.39 1.94 2.98 43.27 30.16
CodeBERT 46.59% 20.36 1.55 2.18 35.16 36.11 49.20% 20.88 1.99 3.14 44.08 30.30
Diversity 39.76% 19.72 1.51 2.14 34.69 33.70 44.40% 20.80 1.99 3.14 44.08 29.97
CAST-F 47.79% 22.04 1.73 2.55 35.79 33.80 50.40% 21.55 2.17 3.38 45.60 31.54
@@
RQ1: How do SOTA LLMs of various scales perform in the code translation task?
@@
In this research question, we evaluate SOTA LLMs across various scales for code translation. As shown in Tab. I, the results reveal several meaningful observations: 1) The LLMs demonstrate impressive performance in the code translation task, highlighting their potential for practical application. Specifically, the proprietary LLMs, such as GPT-4o and GPT-4o-mini, achieve an average CA of 66.37% and 69.72%, respectively, indicating strong potential for LLM-based code translation. 2) A high EM score indicates that the translated code is identical to the ground truth, often used as a metric to assess the extent of training data leakage. Compared to learning-based methods, which ensure no data leakage during training, the EM scores of LLMs are comparable, indicating minimal data leakage. This shows that LLMs can effectively understand the task through prompts and produce semantically correct translated code. 3) Most LLMs outperform learning-based models. Within the same model series, larger models, such as the Gemma2 series, perform better than smaller ones, indicating that scaling up parameters enhances the code translation capabilities of LLMs.
ȷ Finding 1: LLMs demonstrate remarkable performance in code translation, surpassing learning-based methods and showcasing high effectiveness and generalization. Scaling up the parameters of LLMs further enhances their code translation capabilities, highlighting their strong potential for industrial applications.
RQ2: How does CAST perform with SOTA LLMs compared to different ICL exemplar selection strategies?
Referring to the results in Tab. I, we choose three open-source LLMs as primary baselines: Qwen2-7B, the SOTA LLM under 10B parameters; Gemma2-27B, the SOTA LLM under 40B parameters; and Qwen2.5-72B, the SOTA LLM exceeding 70B parameters. CAST-F is used to select exemplars for ICL, and its performance is compared against other exemplar selection strategies. Results in Tab. II demonstrate the effectiveness of CAST-F across LLMs with various parameter scales in the 5-shot code translation task. We conduct t-tests comparing CAST-F with each other ICL retrieval strategies. The resulting p-values indicate that CAST-F significantly outperforms alternative strategies in most settings, with particularly strong statistical advantages observed on GPT-4o and Gemma2-27B. Methods that are highly significantly inferior to our approach ($p<0.05$) are marked with $\bullet$”, and those that are significantly inferior ($0.05\leq p<0.1$) are marked with $\circ$”. There are several key observations: 1) CAST-F consistently outperforms all other ICL exemplar selection strategies across LLMs and translation tasks. Specifically, it surpasses the second-best strategy, Diversity, by 1.6% for Qwen2-7B, 1.9% for Gemma2-27B, and 1.14% for Qwen2.5-72B. This shows that information coverage is a more precise and general measure for ICL, and CAST-F effectively integrates code structural knowledge into LLMs to enhance code translation. Compared with other models, since GPT-4o, as a commercial-closed model, inherently possesses strong capabilities, the performance improvement space brought by structural information injection is not significant, but CAST still maintains a certain advantage compared with baselines. 2) For Gemma2-27B, exemplar selection using LD, CodeBERT, and Diversity performs worse than random sampling. This highlights the importance of effective exemplar selection strategies in ICL. CAST-F, the only list-wise method that considers both information coverage and code structural knowledge, consistently achieves superior results compared to other strategies. 3) AST-ED focuses more on measuring the similarity of the overall structure of the code. In contrast, CAST is based on subtree coverage, placing greater emphasis on matching local structural modules. It provides module-by-module references for translation, effectively reducing translation deviations in local logic and thereby improving the overall performance.
In real-world scenarios, the quality and structure of code repositories vary considerably. To more accurately assess the practical effectiveness of our method, we further evaluated its code translation performance on the AVATAR [35] dataset, which includes 249 Java samples, 250 Python samples, and 6,255 test cases. In addition to computation accuracy, we also computed a range of metrics, including BLEU, Weighted N-Gram (WNG) Match Score, Syntactic Match Score, Dataflow Match Score, and CodeBLEU (weighted combination of the above four parts) [46], offering a more comprehensive evaluation of translation quality. The experimental results, presented in Tab. III, reveal significant performance differences across various ICL strategies in 5-shot code translation using Qwen2-7B on AVATAR. Among the evaluated methods, CAST-F consistently outperforms others in both Java-to-Python and Python-to-Java translation tasks, achieving the highest computational accuracy at 47.79% and 50.40%, respectively, as well as the best overall CodeBLEU scores. These results underscore CAST’s strong generalization ability and its practical effectiveness in real-world code translation tasks. Specifically, CAST-F excels not only in BLEU and WNG but also in Syntactic and dataflow alignment, demonstrating its ability to capture both surface-level and structural aspects of code. In contrast, baseline strategies like Random and Diversity yield lower scores across most metrics, highlighting the importance of informed example selection in ICL-based code translation.
ȷ Finding 2: CAST-F demonstrates robust performance across various LLMs, translation tasks, real-world scenarios, and multiple metrics, emphasizing its ability to capture both surface-level and structural features of code. Consequently, information coverage provides a more precise and generalized measure for ICL, with CAST-F effectively incorporating code structural knowledge into LLMs to improve code translation.
@@
RQ3: How does ICL perform using different exemplar selection strategies as the number of exemplars increases?
Results in Tab. IV and Fig. 3 show the performance of Qwen2-7B in code translation using various ICL exemplar selection strategies, with the number of exemplars increasing from $1$-shot to $20$-shots. Several key observations arise from this analysis: 1) CAST-F consistently achieves the highest average performance across all strategies, with an accuracy of 65.11% at $20$-shots. 2) Increasing the number of exemplars generally improves performance across most strategies. CAST-F demonstrates superior scalability, particularly at higher shot counts, underscoring its robustness in improving code translation. 3) Compared to baselines, CAST-F consistently delivers incremental gains by leveraging information coverage and code structural knowledge. 4) The Diversity strategy ranks second, performing well by considering both point-wise similarity and diversity, but it lacks the ability to account for code structural knowledge or perform selection in a compositional manner. 5) CAST-F achieves the highest accuracy across challenging language pairs, such as Python to Java (58.10%) and Java to Python (69.37%), demonstrating its effectiveness in translating between syntactically diverse languages. 6) A law of diminishing marginal utility is evident in Tab. IV and Fig. 3 as the number of exemplars increases. While performance improves significantly with the first few exemplars, additional gains become minimal beyond $5$-shots due to redundancy or information saturation. However, Qwen2-7B can remit this issue effectively by considering information coverage as a more precise and general measure.
ȷ Finding 3: Despite diminishing marginal utility with more exemplars in all ICL strategies, CAST-F shows superior performance and scalability. It achieves the highest accuracy at higher shot counts by considering information coverage and code structure.
TABLE IV: Performance of Qwen2-7B on code translation with various ICL exemplar selection strategies and numbers of exemplars.
@@
RQ4: How does CAST-A perform when dynamically determining the selection size?
@@
ȷ Finding 4: CAST-A effectively adapts the selection size based on the CAST threshold, outperforming fixed-size strategies while using fewer exemplars. This dynamic adjustment mitigates diminishing returns and improves scalability, showcasing its robustness in handling varying input complexities.
RQ5: How does the CAST convergence of CAST-F compare with that of similarity-only strategy (e.g., LD)?
@@
ȷ Finding 5: CAST-F accelerates the acquisition of necessary knowledge and ensures higher final gains across all tasks. These advantages demonstrate its effectiveness in optimizing exemplar selection for diverse and complex code translation scenarios.
TABLE V: Results of the gain of CAST-F among different sizes of LLMs, consisting of 2B, 9B, and 27B of Gemma2.
@@
RQ6: Does scaling up the model size automatically lead to sufficient code structural knowledge?
As shown in Tab. II, although translation performance improves with larger models, the improvement from explicitly incorporating the code structural knowledge by CAST-F is much more substantial. Additionally, for a fairer comparison, we evaluate CAST-F with $5$-shots on Gemma2-(2B/9B/27B), all of which are from the same family. The results, shown in Tab. V, indicate that the improvement from CAST-F is not significantly related to model scale. This confirms that the additional code structural knowledge from CAST-F is independent of model scale. Therefore, simply scaling up the model or training data does not result in the sufficient emergence of code structural knowledge, highlighting the necessity of explicitly incorporating code syntax.
ȷ Finding 6: Scaling up model size or training data does not lead to the sufficient emergence of code structural knowledge, underscoring the necessity of explicitly considering code syntactic structure in LLMs.
@@
RQ7: What is the hyperparameter sensitivity?